OpenStudy (anonymous):

find dy/dx if y=4(x^2-5x)(9+x^7)

OpenStudy (inkyvoyd):

u=x^2-5x v=9+x^7 d(uv)/dx=uv'+u'v

OpenStudy (inkyvoyd):

remember to multiply by scalar 4 after getting your answer.

OpenStudy (inkyvoyd):

alternative way is to expand out using foil, but imo that's a waste of time

OpenStudy (anonymous):

Use the chain rule 4 is a constant so keep that in front. \[4\frac{ dt }{ dx } = (x^{2} -5x)\frac{ d }{ dx }(9+x^{7}) + (9+x^{7})\frac{ d }{ dx }(x^{2} +5x)\]

OpenStudy (anonymous):

that t is a y.

OpenStudy (anonymous):

thank you so much!

OpenStudy (loser66):

to me, I will do something like